OpenStudy (babynini):

If h(x)=x+sqrt(x) find f^-1(6)

OpenStudy (photon336):

\[h(x) = x+\sqrt{x}\] and \[f^{-1} (6)\] right?

OpenStudy (babynini):

Correct :)

OpenStudy (photon336):

\[f^{-1} = inverse function \]

OpenStudy (photon336):

This is asking you to take the inverse of the function then evaluate it do you know how to do this?

OpenStudy (babynini):

So first we set y=x+sqrt(x) and then solve that for x?

OpenStudy (photon336):

you need to switch the x and y's in the original function and solve for y

OpenStudy (babynini):

Ah ok, let me see.

OpenStudy (babynini):

so we have x=y+ sqrt(y)

OpenStudy (photon336):

then solve for y

OpenStudy (photon336):

Then you can evaluate this for say f^-1(6)

OpenStudy (babynini):

Would a good first step be (x-y)^2 = y and then of course continue solving

OpenStudy (babynini):

\[x^2=(1+2x+y)y\]lala still simplifying xD

OpenStudy (jdoe0001):

well

OpenStudy (jdoe0001):

your "simplification" looks correct :) BUT the "y" will be pervasive on either side though

OpenStudy (jdoe0001):

now lemme hmm graph the original fuhction, one sec

OpenStudy (jdoe0001):

http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiJ4K3NxcnQoeCkiLCJjb2xvciI6IiNFMzBEMEQifSx7InR5cGUiOjAsImVxIjoieC1zcXJ0KHgpIiwiY29sb3IiOiIjMTkxRkNGIn0seyJ0eXBlIjoxMDAwLCJ3aW5kb3ciOlsiLTQuNjQ5OTk5OTk5OTk5OTk5IiwiMTEuNTk5OTk5OTk5OTk5OTk4IiwiLTIuMjk5OTk5OTk5OTk5OTk5NCIsIjcuNyJdfV0- notice if you give a "vertical line test", it won't pass it, meaning, is not a function, just a relation and only one-to-one "functions" have an "inverse function" and one-to-one functions, need to pass the vertical and horizontal line tests this one fails both which as far as I can tell, is why is ending up in a deadend

OpenStudy (babynini):

mm well I have an answer that I SHOULD be able to arrive too. There's just a lot of algebra I guess xD

OpenStudy (jdoe0001):

right... was thinking the same... it should be decouplable just lengthy

OpenStudy (babynini):

\[\frac{ x^2 }{ 1+2x+y}=y\]

OpenStudy (babynini):

Maybe multiply the left side by y?

OpenStudy (jdoe0001):

hmmm

OpenStudy (babynini):

or add y to the left side..might work

OpenStudy (babynini):

ooh you know what we can just do h(x) = x+sqrt(x) = 6 and solve that and then figure it out from there..no?

OpenStudy (babynini):

Because we know the rule for example; f(6)=17 then f^-1(17)=6

OpenStudy (babynini):

\[x+ \sqrt x=6\]\[x=3\] h(3)=6 \[f^{-1}(6)=3\]

OpenStudy (babynini):

*x=4 so f inverse(6)=4